A trading company buys and sells 10000 bottles of pain-balm every year. The company's cost of placing an order of pain-balm is $100. The holding cost per bottle on inventory is $0.30.

To determine the optimum order quantity and inventory cycle time for the pain-balm bottles.

How many orders should be placed each year?



1
Expert's answer
2021-06-17T19:09:25-0400

Ans:-

1)

The optimum order quantity "k" is the ratio between the cost of placing an order "m" and the holding cost "p,"

"k=\frac mp=\frac{100}{0.3}=333" bottles;


inventory cycle time "t" is the ratio between the optimum order quantity "k" and the quantity of the bottles "n,"


"t=\frac kn=\frac{333}{10000}=\frac{1}{30}" year (or "12" days).


2) Number of orders "s" is the inverse of the inventory cycle time "t,"


"s=\frac 1t=30" orders per year.